# A short note on exponential sums, exceptional characters and primes in short intervals

### 20 Jan 2024 v1

@@ Abstract.
This note is an overview of our new results on primes in short intervals $[x-x^{\alpha},x]$ under the assumption of the existence of exceptional Dirichlet characters. We get a small improvement of a 2004 result of Friedlander and Iwaniec. By using Nowak’s or our new estimation of exponential sums, we extend the previous admissible range $0.4937\leqslant\alpha\leqslant 1$ to $0.4925\leqslant\alpha\leqslant 1$.
@@ 1. Introduction
The famous Riemann Hypothesis is equivalent to the asymptotic formula
$\psi(x):=\sum_{n\leqslant x}\Lambda(n)=x+O\left(x^{\frac{1}{2}+\varepsilon}\right),$ (1)
where $\Lambda(n)$ is the von Mangoldt function. This implies that
$\psi(x)-\psi(x-x^{\alpha})=x^{\alpha}+O\left(x^{\frac{1}{2}+\varepsilon}\right)$ (2)
@@
In this paper, we further combine Nowak’s work with Bourgain’s decoupling method (see [2]) and get some similar estimations as Theorem 1.3. We will give a detailed proof of one of our results which is the following one:
@@
$\Delta\ll x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{24851}{142044}}D_{\max}^{\frac{95873}{284088}}x^{\frac{46171}{94696}}+D^{\frac{38017}{144981}}D_{\max}^{\frac{3986}{16109}}x^{\frac{71090}{144981}}+D^{\frac{856679}{2404602}}D_{\max}^{\frac{1281}{6764}}x^{\frac{2185055}{4809204}}\right).$
Obviously, Theorems 1.31.4 imply the same bound $\Delta\ll D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon}$. Then by the similar arguments as in [3], we can get the following result on primes in short intervals:
@@
Let $\chi$ be the real primitive character of conductor D. Let $x\geqslant D^{r}$ with $r=1100000$ and $\frac{2500}{5077}\leqslant\alpha\leqslant 1$. Then
@@
$L(1,\chi)\ll(\log D)^{-1100000^{1100000}-1}$ (4)
if $D$ is a positive power of $x$. In fact, Zhang [9] posited that there is no $L$-function with
$L(1,\chi)<c(\log D)^{-2022}$ (5)
for some constant $c$.
@@ 2. Higher order derivative tests for single exponential sums
where $a_{4}=\frac{1}{2},b_{4}=\frac{32}{205}+\varepsilon,\xi_{4}=0,\eta_{4}=\frac{31}{84}+\varepsilon,\alpha_{4}=\frac{871}{1086},\gamma_{4}=1,\delta_{4}=\frac{1}{2}$, and for every $j>4$,
$\begin{gathered}a_{j}=\frac{a_{j-1}+b_{j-1}+1}{2\left(b_{j-1}+1\right)},\quad b_{j}=\frac{b_{j-1}}{2\left(b_{j-1}+1\right)},\\ \xi_{j}=\frac{\left(\xi_{j-1}+1\right)\left(b_{j-1}+1\right)-a_{j-1}\eta_{j-1}}{2\left(b_{j-1}+1\right)},\quad\eta_{j}=\frac{\eta_{j-1}}{2\left(b_{j-1}+1\right)},\\ \alpha_{j}=\frac{\alpha_{j-1}+1}{2},\\ \gamma_{j}=\frac{a_{j-1}\delta_{j-1}+\left(b_{j-1}+1\right)\left(\gamma_{j-1}+1\right)}{2\left(b_{j-1}+1\right)},\quad\delta_{j}=\frac{\delta_{j-1}}{2\left(b_{j-1}+1\right)}.\end{gathered}$
@@
The proof is very similar to that of [[6], Theorem 1]. The only difference is that we use the bound $E\ll M^{\frac{1}{2}}T^{\frac{13}{84}+\varepsilon}$ in the range $T^{\frac{17}{42}+\varepsilon}\ll M\ll T^{\frac{1}{2}}$. ∎
@@
$E\ll M^{\frac{679}{948}}T^{\frac{16}{237}+\varepsilon}+M^{\frac{33461}{79632}}T^{\frac{6355}{39816}+\varepsilon}+M^{\frac{1957}{2172}}+M^{\frac{2101}{1896}}T^{-\frac{205}{948}}.$
@@ 3. A problem considered by Friedlander, Iwaniec and Nowak
$\ll$ $D_{\max}^{\frac{1}{2}}\left(N_{3}^{\frac{679}{948}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{16}{237}+\varepsilon}+N_{3}^{\frac{33461}{79632}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{6355}{39816}+\varepsilon}+N_{3}^{\frac{1957}{2172}}+N_{3}^{\frac{2101}{1896}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{-\frac{205}{948}}\right).$ (8)
@@
$\ll$ $D^{\frac{1}{6}}x^{\frac{1}{3}}P^{\frac{1}{3}}D_{\max}^{\frac{1}{2}}\left(\frac{P^{\frac{16}{711}}x^{\frac{16}{711}}}{D^{\frac{16}{711}}N_{3}^{\frac{269}{948}}}+\frac{P^{\frac{6355}{119448}}x^{\frac{6355}{119448}}}{D^{\frac{6355}{119448}}N_{3}^{\frac{46171}{79632}}}+N_{3}^{-\frac{215}{2172}}+\frac{D^{\frac{205}{2844}}N_{3}^{\frac{205}{1896}}}{P^{\frac{205}{2844}}x^{\frac{205}{2844}}}\right)$
$\ll$ $D_{\max}^{\frac{1}{2}}\left(\frac{D^{\frac{205}{1422}}P^{\frac{253}{711}}x^{\frac{253}{711}}}{N_{3}^{\frac{269}{948}}}+\frac{D^{\frac{13553}{119448}}P^{\frac{46171}{119448}}x^{\frac{46171}{119448}}}{N_{3}^{\frac{46171}{79632}}}+\frac{D^{\frac{1}{6}}P^{\frac{1}{3}}x^{\frac{1}{3}}}{N_{3}^{\frac{215}{2172}}}+D^{\frac{679}{2844}}N_{3}^{\frac{205}{1896}}P^{\frac{743}{2844}}x^{\frac{743}{2844}}\right).$ (9)
@@
$\Delta\ll$ $\left(\frac{Dx^{2}}{N}\right)^{\frac{1}{3}}+D_{\max}^{\frac{1}{2}}\left(D^{\frac{205}{1422}}N^{\frac{743}{2844}}x^{\frac{253}{711}}+D^{\frac{13553}{119448}}N^{\frac{46171}{238896}}x^{\frac{46171}{119448}}+D^{\frac{1}{6}}N^{\frac{1957}{6516}}x^{\frac{1}{3}}+D^{\frac{679}{2844}}N^{\frac{2101}{5688}}x^{\frac{743}{2844}}\right).$ (10)
@@
$N=D^{\frac{538}{1691}}D_{\max}^{-\frac{1422}{1691}}x^{\frac{884}{1691}},$
@@
$\Delta\ll$ $x^{\varepsilon}\left(D^{\frac{1153}{5073}}D_{\max}^{\frac{474}{1691}}x^{\frac{2498}{5073}}+D^{\frac{24851}{142044}}D_{\max}^{\frac{95873}{284088}}x^{\frac{46171}{94696}}+D^{\frac{38017}{144981}}D_{\max}^{\frac{3986}{16109}}x^{\frac{71090}{144981}}+D^{\frac{856679}{2404602}}D_{\max}^{\frac{1281}{6764}}x^{\frac{2185055}{4809204}}\right)$
$\ll$ $D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon},$ (11)
which is our Theorem 1.4, implies the same bound as Nowak’s Theorem 1.3(A) but reduces the exponent of $x$ on the second term on the right hand side (we have $0.48990.4898<\frac{26507}{54112}<$ 0.4899 and $0.48760.4875<\frac{46171}{94696}<$ 0.4876).
@@ 4. A new application: Primes in Short Intervals
In his articles [6][8], Nowak considered several problems concerning products of $L$-series and pointed that his method is more powerful in the case of exponential sums depending on several parameters. For exponential sums depending on one parameter only, multiple sum estimations are often more powerful. (see [[8], Concluding Remark]). Now we follow Friedlander and Iwaniec directly and show that our Theorem 1.4 can be used to extend the range of $\alpha$. We first introduce some arithmetic functions:
@@
$\mathcal{D}\left(x;\rho\right)=$ $L(1,\chi)x\log x+\left(L^{\prime}(1,\chi)+(2\gamma-1)L(1,\chi)\right)x+O\left(D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon}\right),$ (14)
@@
These can be proved by similar arguments as in [4], and the only difference is that we use our Theorem 1.4 instead of Theorem 1.2 in order to prove the formula (14). Actually (14) is the Corollary 1.1 in [7]. ∎
@@
By Lemma 4.1 and similar arguments as in [3], for $D^{2}x^{\frac{2499}{5075}}<y\leqslant x$ we obtain the following asymptotic formula
@@
For $D^{\frac{5}{2}}x^{\frac{2499}{5075}+\frac{2576}{5075r}}<y\leqslant x$ with $r\geqslant 1$, we have
@@
subject to the conditions of Lemma 4.3. We require $D\leqslant x^{\frac{1}{r}}$ and then choose a number $\theta$ with $\frac{2499}{5075}<\theta<\frac{1}{2}$ satisfies
$\frac{5}{2r}+\frac{2499}{5075}+\frac{2576}{5075r}\leqslant\theta.$
By choosing $\theta=\frac{2500}{5077}$ and $r=1100000$, we complete the proof of Theorem 1.5. We remark that if we only use Bourgain’s result but without using Huxley’s one, we can only obtain the range $0.4931\leqslant\alpha\leqslant 1$. It seems that if we are able to make some modifications, our result may be improved to $0.4923\leqslant\alpha\leqslant 1$, which is rather near to the limit obtained by this method.

### 30 Jan 2024 v2

@@ Abstract.
We focus on primes in short intervals $[x-x^{\alpha},x]$ under the assumption of the existence of exceptional Dirichlet characters and get a small improvement of a 2004 result of Friedlander and Iwaniec. By using our new estimation of exponential sums, we extend the previous admissible range $0.4937\leqslant\alpha\leqslant 1$ to $0.4923\leqslant\alpha\leqslant 1$.
@@ 1. Introduction
The famous Riemann Hypothesis is equivalent to the asymptotic formulas
$\psi(x):=\sum_{n\leqslant x}\Lambda(n)=x+O\left(x^{\frac{1}{2}+\varepsilon}\right)\quad\operatorname{and}\quad\pi(x)=\operatorname{Li}(x)+O\left(x^{\frac{1}{2}+\varepsilon}\right),$ (1)
where $\Lambda(n)$ is the von Mangoldt function. These imply that
$\psi(x)-\psi(x-x^{\alpha})=x^{\alpha}+O\left(x^{\frac{1}{2}+\varepsilon}\right)\quad\operatorname{and}\quad\pi(x)-\pi(x-x^{\alpha})=\left(\operatorname{Li}(x)-\operatorname{Li}(x-x^{\alpha})\right)+O\left(x^{\frac{1}{2}+\varepsilon}\right)$ (2)
@@
and
$\pi(x)-\pi(x-x^{\alpha})=\left(\operatorname{Li}(x)-\operatorname{Li}(x-x^{\alpha})\right)\left\{1+O\left(L(1,\chi)(\log x)^{r^{r}}\right)\right\},$
@@
Obviously, both (A) and (B) in Theorem 1.3 imply the same bound $\Delta\ll D^{\frac{2575}{5073}}x^{\frac{2498}{5073}+\varepsilon}$. In this paper, we further combine Nowak’s work with Bourgain’s new bound (see [2]) and get some improvement on Theorem 1.3. We will give a detailed proof of one of our results which is the following one:
@@
$\Delta\ll$ $x^{\varepsilon}\left(D^{\frac{118}{519}}D_{\max}^{\frac{97}{346}}x^{\frac{511}{1038}}+D^{\frac{121}{692}}D_{\max}^{\frac{467}{1384}}x^{\frac{675}{1384}}+D^{\frac{56039}{213309}}D_{\max}^{\frac{69941}{284412}}x^{\frac{419257}{853236}}+D^{\frac{17936}{50343}}D_{\max}^{\frac{131}{692}}x^{\frac{91507}{201372}}\right)$
$\ll$ $D^{\frac{527}{1038}}x^{\frac{511}{1038}+\varepsilon}.$
Then by the similar arguments as in [3], we can get the following result on primes in short intervals:
@@
Let $\chi$ be the real primitive character of conductor D. Let $x\geqslant D^{r}$ with $r=433433$ and $0.4923\leqslant\alpha\leqslant 1$. Then
@@
and
$\pi(x)-\pi(x-x^{\alpha})=\left(\operatorname{Li}(x)-\operatorname{Li}(x-x^{\alpha})\right)\left\{1+O\left(L(1,\chi)(\log x)^{r^{r}}\right)\right\}.$
@@
$L(1,\chi)\ll(\log D)^{-433433^{433433}-1}$ (4)
if $D$ is a positive power of $x$. Unfortunately, Zhang [10] posited that there is no $L$-function with
$L(1,\chi)\ll(\log D)^{-2022}.$ (5)
@@ 2. Higher order derivative tests for single exponential sums
where $a_{4}=\frac{1}{2},b_{4}=\frac{13}{84}+\varepsilon,\xi_{4}=0,\eta_{4}=\frac{31}{84}+\varepsilon,\alpha_{4}=\frac{334}{411},\gamma_{4}=1,\delta_{4}=\frac{1}{2}$, and for every $j>4$,
$\begin{gathered}a_{j}=\frac{a_{j-1}+b_{j-1}+1}{2\left(b_{j-1}+1\right)},\quad b_{j}=\frac{b_{j-1}}{2\left(b_{j-1}+1\right)},\\ \xi_{j}=\frac{\left(\xi_{j-1}+1\right)\left(b_{j-1}+1\right)-a_{j-1}\eta_{j-1}}{2\left(b_{j-1}+1\right)},\quad\eta_{j}=\frac{\eta_{j-1}}{2\left(b_{j-1}+1\right)},\quad\alpha_{j}=\frac{\alpha_{j-1}+1}{2},\\ \gamma_{j}=\frac{a_{j-1}\delta_{j-1}+\left(b_{j-1}+1\right)\left(\gamma_{j-1}+1\right)}{2\left(b_{j-1}+1\right)},\quad\delta_{j}=\frac{\delta_{j-1}}{2\left(b_{j-1}+1\right)}.\end{gathered}$
@@
The proof is very similar to that of [[7], Theorem 1]. The only difference is that we use the bound $E\ll M^{\frac{1}{2}}T^{\frac{13}{84}+\varepsilon}$ in the range $T^{\frac{17}{42}+\varepsilon}\ll M\ll T^{\frac{1}{2}}$ and the bound $E\ll M^{\frac{1}{2}}T^{\frac{32}{205}+\varepsilon}\ll M^{\frac{333}{410}+\varepsilon}\ll M^{\frac{334}{411}}$ in the range $T^{\frac{1}{2}}\ll M\ll T^{\frac{181}{328}+\varepsilon}$. ∎
@@
$E\ll M^{\frac{139}{194}}T^{\frac{13}{194}+\varepsilon}+M^{\frac{163}{388}}T^{\frac{31}{194}+\varepsilon}+M^{\frac{745}{822}}+M^{\frac{215}{194}}T^{-\frac{21}{97}}.$
@@ 3. A problem considered by Friedlander, Iwaniec and Nowak
$\ll$ $D_{\max}^{\frac{1}{2}}\left(N_{3}^{\frac{139}{194}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{13}{194}+\varepsilon}+N_{3}^{\frac{163}{388}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{\frac{31}{194}+\varepsilon}+N_{3}^{\frac{745}{822}}+N_{3}^{\frac{215}{194}}\left(\left(\frac{Px}{D}\right)^{\frac{1}{3}}\right)^{-\frac{21}{97}}\right).$ (8)
@@
$\ll$ $D^{\frac{1}{6}}x^{\frac{1}{3}}P^{\frac{1}{3}}D_{\max}^{\frac{1}{2}}\left(D^{-\frac{13}{582}}N_{3}^{-\frac{55}{194}}P^{\frac{13}{582}}x^{\frac{13}{582}}+D^{-\frac{31}{582}}N_{3}^{-\frac{225}{388}}P^{\frac{31}{582}}x^{\frac{31}{582}}+N_{3}^{-\frac{77}{822}}+D^{\frac{7}{97}}N_{3}^{\frac{21}{194}}P^{-\frac{7}{97}}x^{-\frac{7}{97}}\right)$
$\ll$ $D_{\max}^{\frac{1}{2}}\left(D^{\frac{14}{97}}N_{3}^{-\frac{55}{194}}P^{\frac{69}{194}}x^{\frac{69}{194}}+D^{\frac{11}{97}}N_{3}^{-\frac{225}{388}}P^{\frac{75}{194}}x^{\frac{75}{194}}+D^{\frac{1}{6}}N_{3}^{-\frac{77}{822}}P^{\frac{1}{3}}x^{\frac{1}{3}}+D^{\frac{139}{582}}N_{3}^{\frac{21}{194}}P^{\frac{76}{291}}x^{\frac{76}{291}}\right).$ (9)
@@
$\Delta\ll$ $\left(\frac{Dx^{2}}{N}\right)^{\frac{1}{3}}+D_{\max}^{\frac{1}{2}}\left(D^{\frac{14}{97}}N^{\frac{76}{291}}x^{\frac{69}{194}}+D^{\frac{11}{97}}N^{\frac{75}{388}}x^{\frac{75}{194}}+D^{\frac{1}{6}}N^{\frac{745}{2466}}x^{\frac{1}{3}}+D^{\frac{139}{582}}N^{\frac{215}{582}}x^{\frac{76}{291}}\right).$ (10)
@@
$N=D^{\frac{55}{173}}D_{\max}^{-\frac{291}{346}}x^{\frac{181}{346}},$
@@
$\Delta\ll$ $x^{\varepsilon}\left(D^{\frac{118}{519}}D_{\max}^{\frac{97}{346}}x^{\frac{511}{1038}}+D^{\frac{121}{692}}D_{\max}^{\frac{467}{1384}}x^{\frac{675}{1384}}+D^{\frac{56039}{213309}}D_{\max}^{\frac{69941}{284412}}x^{\frac{419257}{853236}}+D^{\frac{17936}{50343}}D_{\max}^{\frac{131}{692}}x^{\frac{91507}{201372}}\right)$
$\ll$ $D^{\frac{527}{1038}}x^{\frac{511}{1038}+\varepsilon},$ (11)
which is our Theorem 1.4, improves Nowak’s Theorem 1.3(A) (we have $0.49250.4924<\frac{2498}{5073}<$ 0.4925 and $0.49230.4922<\frac{511}{1038}<$ 0.4923).
@@ 4. A new application: Primes in Short Intervals
In his articles [7][9], Nowak considered several problems concerning products of $L$-series and pointed that his method is more powerful in the case of exponential sums depending on several parameters. For exponential sums depending on one parameter only, multiple sum estimations are often more powerful. (see [[9], Concluding Remark]). Clearly our new estimation can improve these results. Now we follow Friedlander and Iwaniec directly and show that our Theorem 1.4 can be used to extend the range of $\alpha$. We first introduce some arithmetic functions:
@@
$\mathcal{D}\left(x;\rho\right)=$ $L(1,\chi)x\log x+\left(L^{\prime}(1,\chi)+(2\gamma-1)L(1,\chi)\right)x+O\left(D^{\frac{527}{1038}}x^{\frac{511}{1038}+\varepsilon}\right),$ (14)
@@
These can be proved by similar arguments as in [4], and the only difference is that we use our Theorem 1.4 instead of Theorem 1.2 in order to prove the formula (14). ∎
@@
By Lemma 4.1 and similar arguments as in [3], for $D^{2}x^{\frac{492293}{1000000}}<y\leqslant x$ we obtain the following asymptotic formula
@@
For $D^{\frac{5}{2}}x^{\frac{492293}{1000000}+\frac{507707}{1000000r}}<y\leqslant x$ with $r\geqslant 1$, we have
@@
subject to the conditions of Lemma 4.3. We require $D\leqslant x^{\frac{1}{r}}$ and then choose a number $\theta$ with $\frac{492293}{1000000}<\theta<\frac{1}{2}$ satisfies
$\frac{5}{2r}+\frac{492293}{1000000}+\frac{507707}{1000000r}\leqslant\theta.$
By choosing $\theta=0.4923$ and $r=433433$, we complete the proof of Theorem 1.5. We remark that the range $0.4923\leqslant\alpha\leqslant 1$ is rather near to the limit obtained by this method. In his preprint [6], Merikoski mentioned that some sieve arguments can be used to this problem.